Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
amagazzu - 10:07 Saturday 15 May 2021 (51799) Print this report
Operator Report - Initial Status

ITF State: Infrared cavities unlocked
ITF Mode: Commissioning.
Quick Summary: IMC and RFC locked; all Suspensions loops closed, SR IP data missing; all SBEs loops closed, SDB2 tracking loop Open; NEB_Eurotherm device not responding.
Activities ongoing since this morning: Preparation work for tests on IMC instrumented baffle (Menendez, Karathanasis, Chiummo);

Images attached to this report
Comments to this report:
Boschi - 15:48 Saturday 15 May 2021 (51802) Print this report

SR Data are now back. Top crate has been rebooted. All controls are now closed.

Search Help